OpenStudy (tmostiller):

What is the solution of the system? Use elimination. -12x-y=6 17x+y=4

OpenStudy (mathmate):

Elimination means that you add or subtract multiples of the equations so that one of the variables cancel out. Here, noting that -y appears in the first equation, and +y appears in the second. So by simply adding the first equation to the second, we will have eliminated y, resulting in an equation containing x only. Solve this last equation for x, and back substitute in the first equation to find y. Finally, using the calculated values of x and y, substitute back into EACH of the equations to make sure that you have the correct answer. Example: 4x+3y=7 2x-3y=-1 Here again, we have +3y and -3y, we simply add to eliminate y. (4x+2x)+0y=7-1 6x=6 x=1 Put x=1 into the first equation to solve for y 4(1)+3y=7 3y=7-4 3y=3 y=1 Now put x=1, y=1 into the first equation: 4(1)+3(1)=4+3=7 ....checks 2(1)-3(1)=2-3=-1 .... checks So (x,y)=(1,1)

OpenStudy (tmostiller):

Thank you